# Measuring Casimir Force Across a Superconducting Transition

### v1

@@ VI.2 Experimentally Isolating Casimir Forces
To isolate the Casimir force in our solid-state parallel-plate cavities, we systematically eliminate other forces acting on the superconducting membrane, including electrostatic attractions, van der Waals (VdW) interactions, and thermal-elastic effects from temperature sweeps across $T_{C}$. Our protocol begins by canceling electrostatic forces between the membrane and backgate, followed by tip-membrane electrostatics, and finally minimizing VdW interactions. Combined with precise STM positioning and detailed temperature control, this sequence enables reproducible, high-sensitivity measurements of Casimir forces across the superconducting transition.
VI.3 Positioning and Minimizing Tip-Membrane Forces
@@
As shown in Fig. 5a, the STM’s enables to scan the surface of the membrane within a $3\times 3$ nm2 area, effectively imaging the mechanical resonator at the scale of individual atoms. The surface lacks crystalline smoothness due to sputtering growth. The resulting roughness of 1-2 nm results in high and low points in the scan window. By combining the STM’s current-to-tip height feedback system with a local gradient search algorithm, the STM autonomously locks onto the locally highest point with 10 pm precision — nearly 30 times smaller than the 300 pm radius of a niobium atom on the NbTiN membrane (see inset Fig. 5a). This remarkable sub-atomic spatial resolution ensures the tip remains fixed at the same membrane location throughout measurements, minimizing positional drift and enhancing reproducibility. This level of precision ensures that the local force environment remains stable and consistent throughout measurements. While tunneling current depends exponentially on tip-sample separation, simply maintaining a constant current at a different location would not guarantee identical forces. By fixing the tip at the same membrane position with sub-atomic precision, we minimize variations in local interactions, ensuring reproducible force measurements. With this sub-atomic localization, the STM provides extreme control over how our measurements interact with the experimental system. This control is essential for precisely canceling out any perturbative effects introduced by the measurement process.
VI.4 Mitigating Electrostatic Forces
@@
To minimize electrostatic forces between the STM tip and the membrane, the tip–membrane bias voltage is swept while maintaining constant tip–sample separation. The resonance frequency response is monitored to identify the compensation voltage where residual interactions vanish (Fig. 5d). This procedure ensures that tip-induced electrostatic forces remain negligible throughout measurements.
VI.5 Tip-membrane van der Waals forces
As the STM tip approaches the membrane, VdW forces between the tip and membrane can become significant. To mitigate this, the STM tip is sharpened and cleaned by dipping it into a copper sample and applying voltage pulses. Atomic steps on the copper surface are imaged to verify sharpness, ensuring the tip’s outermost atoms are well-defined. While this method confirms the presence of a clean, localized apex, it does not fully characterize the sharpness beyond this range, as additional atomic-scale features — such as dangling atoms — may still be present. The optimal landing spot on the membrane is then identified as the highest, most stable location to minimize position-dependent van der Waals interactions during temperature sweeps. The membrane’s resonance frequency response is validated at different tip-membrane currents (i.e. different tip-membrane distances) to ensure Lorentzian mechanical peaks, confirming that van der Waals forces are minimized (Fig. 5e); the lack of frequency shift across these distances indicates that distance-dependent VdW contributions are negligible within our measurement uncertainty.
VI.6 Managing Thermal Drift and Elastic Forces
@@
Temperature sweeps introduce thermal drift and elastic effects that can distort measurements if not properly managed. To address this, we begin at the base temperature (4.45 K) with a precise $3\times 3$ nm2 scan of the membrane. Drift is actively corrected using the STM’s scanning capabilities at atomic scales and feedback system operating at frequencies up to 50 Hz, maintaining a lateral tip position to within a 10 pm region during resonance sweeps.
As the temperature increases, thermal expansion induces additional drift, which is mitigated by incremental heating and stabilization protocols. For temperature spans of 0.1 K, 1 K, and 5 K, stabilization times of 1 hour, 2.5 hours, and 3 hours, respectively, are employed to maintain precise alignment. These steps ensure that thermal-elastic effects are systematically accounted for during temperature-dependent measurements.
Following the compensation of electrostatic forces and van der Waals and forces, the system is prepared for long-range temperature sweeps to study Casimir force variations from a base temperature of 4.45 K to the superconducting transition temperature, $T_{C}\sim 14.2$ K. These large sweeps provide an overall understanding of how the forces between the plates evolve over a broad temperature range, albeit with reduced stability compared to small-range sweeps near $T_{C}$. The minimized residual interactions, combined with the STM’s high sensitivity, establish a robust platform for detecting subtle changes in the Casimir force.
However, measuring Casimir force differences over large temperature spans introduces additional challenges. These include variability in the mechanical properties of the NbTiN membranes, shifts in the tip landing position, and increased thermalization time requirements, all of which can influence interaction consistency. For this reason, the most critical measurements are conducted over small temperature ranges near the superconducting transition, where the interplay between superconductivity and the Casimir force is perhaps more prominent. As illustrated in Fig. 5f, large sweeps serve as a broader diagnostic tool, while precision measurements across $T_{C}$ provide the most meaningful insights into the behavior of the Casimir force in superconducting systems.
For comparison, Supplementary Information C includes cases where the tip-membrane interaction is moderately larger or significantly stronger. These examples highlight the critical importance of minimizing VdW and electrostatic forces to ensure reliable Casimir force measurements. All of these procedures are performed for every experimental run across $T_{C}$. By achieving these conditions, our experimental setup is optimized for detecting subtle variations in the Casimir force.
@@ Acknowledgements.
Quality factor $Q$ ($\sim 4.45$ K) (×$\times 10^{3}$) 720 200 - 300
Quality factor $Q$ ($\sim 14.2$ K) (×$\times 10^{3}$) 1000 130 - 170
@@
    Wafer Preparation: A double-side polished 300 $\mu$m bare Si wafer was selected. This thickness was chosen to avoid fragility associated with wafers thinner than 200 $\mu$m, while minimizing the measurement uncertainty of thin-film stresses, which increases in wafers thicker than 500 $\mu$m.
@@
    Insulating Layer Deposition: A 1 $\mu$m thick layer of low pressure chemical vapor deposition (LPCVD) wet thermal oxide (SiO2) was deposited on the Si wafer to form a robust insulating barrier.
@@
    Cryogenic SF6 Undercut: The final undercut is performed at –120 C using SF6 plasma. Prior to etching, the sample is cooled down in a high-vacuum environment with a 10 sccm helium flow for 15 min. The etch is then conducted for 3 min and 10 s with SF6 plasma, where the RF power is set to 0 W and the ICP power to 3000 W. After etching, the sample is warmed to 40 C, pumped out, and then unloaded from the process chamber.
@@
    High-Vacuum Baking: The STM holder is placed inside the STM load lock and baked at 200 C under high vacuum ($<10^{-8}$ mbar) for two days. This step removes contaminants such as native oxides and residual organic materials.
@@
    Tip Sharpening on an Atomically Flat Copper Sample: Prior to measuring the Casimir samples, an atomically flat copper sample is used to pre-condition the STM tip. This involves: (1) Pulsing the tip with appropriate voltages, and (2) Dipping the tip into the copper sample. These steps refine the tip’s profile, yielding atomic resolution as confirmed by imaging a step height portfolio on the copper sample, as shown in Fig. S10(b).
@@
    Positioning the STM Tip on Casimir Cavities: The in-situ optical camera is employed for recognition of the membrane’s location, then the STM tip can be positioned on the spot, where is approximately 50 $\mu$m from the edge of the Casimir cavity. This placement minimizes unwanted interactions between the tip and the suspended membrane. Once positioned, the tip is gently dipped into the membrane by approximately 2 nm for a few seconds. This ensures that the measurement spot conforms well to the apex of the tip, facilitating high-resolution scans over a $100\times 100$ nm2 area. When the STM tip can perform a high-resolution scan over $100\times 100~{}\rm nm^{2}$ around a chosen spot, we consider the chosen spot to be stable enough for measurements.
@@
    Temperature Sweep Considerations: During measurements of the membrane’s resonance frequency as a function of temperature, it is important that the STM tip cannot be retracted out of the tunneling range in the $z$-direction, since the piezo-actuator’s expansion or contraction in the $x$/$y$-directions with temperature variations can cause the loss of the original measurement spot. As a result, the atomic tracker, a feedback system in the STM for tracking the same spot on the topography, is activated not only during high-resolution measurement (with a radius of 10 pm) but also during temperature changes (with a tracking radius of approximately 100 pm).
@@
$\Delta\omega^{2}=-\frac{\epsilon_{0}}{\rho t}\frac{(V_{\rm bg}-V_{0})^{2}}{d^{3}}-\frac{\epsilon_{0}}{\rho t}\frac{\delta V_{\rm bg}^{2}}{d^{3}}~{},$ (S.1)
where $\epsilon_{0}$ is the vacuum permittivity, $\rho$ is the film density, $t$ is the membrane thickness, $V_{\rm bg}$ is the backgate voltage, and $V_{0}$ is the residual electrostatic voltage. Fitting the experimental data with the above formula, yields the densities ($\rho$) of the suspended NbTiN membranes, while the peak frequencies of the parabolic response curves provide the film stresses.
Supplementary Information C4: Spectroscopy ($I$-$V$) measurement for checking the bias voltage’s influence on the membrane frequency
Figure S12: (a) Schematic of the spectroscopy ($I$-$V$) measurement, where the separation between the STM tip and the suspended membrane is kept the same, by varying the bias voltage between the STM and the membrane, and turning off the feedback control of the tunneling current. Tunneling current $I$ as a function of the bias voltage ($V$), between the STM tip and the (b) small- and (c) big-gap membranes, for each pair of $I$-$V$ values, the tip-sample separation are kept. These values are used for checking whether the tip-sample interaction depends on the bias voltage, as well as the dependence on the temperature, as shown in (d) for small-gap membrane, and in (e) for big-gap membrane.
Experimental results on the nanomembranes’ resonant frequencies dependence on the bias voltage is shown in Fig. 3(d) in the main text. The experiment is performed by setting the separation between the STM tip and the membrane unchanged, and varies the bias voltage and thus the tunneling current between the tip and the membrane. The corresponding tunneling current values at different voltages are shown in Fig. S12(b) for small-gap and (c) for big-gap membranes. The insets in Fig. 3(d) are shown with more details in Fig. S12(d) for small-gap and (e) for big-gap membranes
Supplementary Information D: Measurements with atomic force microscopy
@@
We performed atomic force microscopy (AFM) measurement to identify the roughness of the NbTiN suspended membranes and backgates (bottom films) after the cryogenic SF6 plasma undercut. In Fig. S13, we show the AFM scan of the backgate films on the small-gap (Panel (a)) and big-gap (Panel (c)) chips, whose RMS roughness are $(0.90\pm 0.05)$ nm and $(0.96\pm 0.07)$ nm, respectively. The area exposed to the plasma etchant is roughened up to $(1.73\pm 0.55)$ nm and $(1.78\pm 0.80)$ nm for the small- and big-gap bottom films, respectively. The zoom-in images of the bottom films under the small- and big-gap cavities, are shown in Panel (b) and Panel (d), respectively, showing the patch sizes of the NbTiN films to be around $\ell=20\sim 30$ nm in radius.
@@
In Fig. S14, we scan the top of the membranes on small- (Panel (a)) and big-gaps (Panel (b)), after they collapsed to the bottom. The roughness of the top membrane over the small-gap cavity is $(1.50\pm 0.27)$ nm, while the one over the big-gap cavity is $(2.73\pm 0.26)$ nm, from which we know that depositing a thicker amorphous silicon layer (1213 nm instead of 190 nm), will increase the roughness of the top NbTiN film a bit.
@@
By comparing the heights of the collapsed membranes over the small-gap (Panel (b)) and big-gap (Panel (d)) cavities in Fig. S15, we found both the thicknesses of the membranes over the small- and big-gap to be around 155 nm. From Panels (a) and (c), the small-gap size is found to be (345-155)=190 nm, and the big gap size is found to be (1368-155)=1213 nm, both matching the values we measured with the SEM. Here the suspended films are fabricated with large opening windows on the top NbTiN films, thus large overhangs are realized and the film stresses are released. for the convenience of AFM scan,
@@
[…]

### v2

@@ V.3 Experimentally isolating Casimir forces
To isolate the Casimir force in our solid-state parallel-plate cavities, we systematically eliminate other forces acting on the superconducting membrane, including membrane–backgate electrostatics, tip–membrane electrostatics, van der Waals (VdW) interactions, and thermal–elastic effects from temperature sweeps across $T_{C}$. Our protocol cancels membrane–backgate electrostatics first, then tip–membrane electrostatics, and finally minimizes VdW interactions. With the STM tip fixed at a single membrane location and temperature carefully controlled, this sequence enables reproducible, high-sensitivity measurements of Casimir forces across the superconducting transition.
V.3.1 Positioning and Minimizing Tip–Membrane Forces
@@
As shown in Fig. 4b, the STM can scan a $3\times 3$ nm2 area of the membrane, imaging the resonator with atomic resolution. This capability is unique in that the same instrument both resolves the membrane’s resonant motion and provides in-situ atomic-scale images of the exact spot being measured. Due to sputtering growth, the surface has a 1–2 nm roughness, producing local high and low points. Using the STM’s current-to-height feedback and a local gradient search algorithm, the system autonomously locks onto the highest point with 10 pm precision, nearly 30 times smaller than the 300 pm radius of a niobium atom on the NbTiN membrane (inset Fig. 4b). This sub-atomic spatial precision ensures the tip remains fixed at one location, minimizing positional drift and keeping the local force environment stable. Holding the same position is essential since identical current at a different location would not guarantee the same forces [43]. Fixing the tip with sub-atomic precision minimizes variations in local interactions and allows precise cancellation of perturbative effects introduced by the measurement itself.
V.3.2 Mitigating Electrostatic Forces
@@
To ensure minimal electrostatic forces between the STM tip and the membrane, the tip–membrane bias voltage is swept while maintaining constant tip–sample separation. Here we clearly show that irrespective of the compensation voltage between tip and membrane, the frequency shift will be small compared to the measured signal changes in Fig. 4g. This procedure ensures that tip-induced electrostatic forces remain negligible throughout measurements.
V.3.3 Minimizing Tip–Membrane van der Waals Forces
As the STM tip approaches the membrane, van der Waals (VdW) forces can become significant. To reduce these effects, the tip is sharpened and cleaned by dipping into a copper sample and applying voltage pulses. Imaging atomic steps on the copper surface verifies a well-defined apex, although sub-apex features such as dangling atoms may still remain.
The optimal landing site on the membrane is chosen as the highest and most stable location, minimizing position-dependent VdW interactions during temperature sweeps. The membrane’s resonance frequency is then measured at different tip–membrane currents (i.e., separations) to confirm Lorentzian mechanical peaks. As shown in Fig. 4e, the absence of frequency shifts across these distances indicates that distance-dependent VdW contributions are negligible within our measurement uncertainty.
V.3.4 Managing Thermal Drift and Determining Casimir Pressure Gradient Variations
@@
Temperature sweeps introduce both thermal drift and elastic effects that can distort Casimir measurements if not carefully managed. We begin each run at the base temperature (4.45 K) with a precise $3\times 3$ nm2 scan of the membrane, using the STM’s atomic-scale scanning and 50 Hz feedback to maintain the tip position within a 10 pm region during resonance sweeps. Incremental heating with stabilization periods (1 hour for 0.1 K, 2.5 hours for 1 K, and 3 hours for 5 K) minimizes drift from thermal expansion.
Long-range sweeps from 4.45 K to $T_{C}\sim 14.2$ K are performed with open-flow helium cooling (0.45 L/hour) to eliminate vibrations that would otherwise degrade STM performance. The STM holds 9 L of liquid helium — 1.8 L for overnight tip and temperature stabilization and 7.2 L for the $T_{C}$ sweep. Ultra-high vacuum ($\sim 10^{-10}$ mbar) preserves clean surfaces after annealing and suppresses squeezed-film damping at nanoscopic separations.
After systematically eliminating electrostatic forces (membrane–backgate and tip–membrane) and minimizing van der Waals forces, we track resonance frequency shifts across $T_{C}$. Small-gap (190 nm) and big-gap (1213 nm) membranes are measured and placed side-by-side in the STM holder; the large-gap device serves as a reference to remove non-Casimir effects from material property changes, thermal expansion, or tunneling variations. For consistency, we denote temperature-dependent variations by $\Delta\psi$ and fixed-temperature variations by $\delta\psi$ for any quantity $\psi$ (e.g., frequency, pressure).
Independent resistivity sweeps confirm $T_{C}\sim 14.2$ K for both devices (Fig. 4f). This resistance includes non-superconducting leads and thus does not go to zero, but allows one to observe where the abrupt transition happens. Following calibration in Supplementary Information G, we linearly fit the elastic force below $T_{C}$ to extract the residual $\Delta\omega^{2}$ above $T_{C}$, where $\omega=2\pi f$. Because gap size is the only difference between the membranes, differences in $\Delta\omega^{2}$ are attributed to a separation-dependent force. With other distance-dependent forces removed, this shift is assigned to the Casimir force. Using $\Delta\omega^{2}$ from the big-gap device, we estimate mechanical changes in NbTiN and compute the Casimir-induced $\omega^{2}$ shift in the small-gap membrane.
V.4 Converting Mechanical Frequency Change to Pressure Gradient
The relation $\Delta\omega^{2}=-\Delta P^{\prime}(d)/\rho t$ links frequency shift to the pressure gradient, where $d$ is separation, $\rho$ density, and $t$ thickness (Supplementary Information F). From the data, we obtain a Casimir pressure gradient variation of $12.10\pm 1.38$ kPa/m for a $0.20\pm 0.02$ K temperature change (Fig. 4c). For our parallel-plate geometry, $\Delta f/f\approx-\Delta P^{\prime}_{\rm Cas}L^{2}/(4\pi^{2}\sigma t)=-\Delta F^{\prime}_{\rm Cas}/(4\pi^{2}\sigma t)$, where $\sigma$ is stress (Supplementary Information F).
Based on Lifshitz theory [14], the Casimir pressure at $d=190$ nm is
$P_{\rm Cas}|_{d=190\ \mathrm{nm}}=-1.081\times 10^{-24}/d^{3.507}\ \mathrm{[Pa]}\approx-0.403\ \mathrm{Pa}$
($d$ in meters; negative sign indicates attraction). Finite element simulations then yield a Casimir force change $\Delta F_{\rm Cas}=-0.33\pm 0.04$ nN, or $\Delta P_{\rm Cas}=-0.65\pm 0.07$ mPa across $T_{C}$. The negative signs indicate increased attraction above $T_{C}$.
The abrupt pressure gradient change in Fig. 4h is not predicted by existing theory [14] which focuses on long-term drifts below $T_{C}$. The Casimir force, as the derivative of the Helmholtz free energy $F_{\rm Cas}=-\partial{\cal F}/\partial d$ [44], should remain continuous through a second-order phase transition. However, our data agree only with a hybrid plasma-BCS model, which combines plasma permittivity above $T_{C}$ with BCS below $T_{C}$, altering the zero-frequency ($n=0$) Matsubara term. Only this combination produces a discontinuous pressure gradient quantitatively consistent with our measured signal in both direction and order-of-magnitude.
[…]
VI Theoretical Expectations and Sensitivity Targets
Model Above $T_{C}$ Below $T_{C}$ Trend, $\Delta P^{\prime}$ (kPa/m)
Plasma–BCS Plasma BCS jump, 6.0
Plasma–Plasma Plasma Plasma smooth, 0.065
Drude–BCS Drude BCS kink, 0.019
Table 1: Variation of Casimir pressure gradient $\Delta P^{\prime}$ values correspond to changes across $T_{C}$ ($\pm 0.1~\rm K$).
Understanding how superconductivity influences the Casimir effect could open new directions in condensed matter physics [31, 32] and superconductivity [34, 35], and may eventually inform more speculative proposals [38, 39, 40, 41, 42]. Before such questions can be meaningfully addressed, experiments must first reach the sensitivity required to resolve changes predicted by conventional electrodynamic models.
Two baseline targets already impose stringent requirements. Over a broad temperature sweep from $T_{C}\approx 14~\mathrm{K}$ down to a few kelvin, the predicted difference between plasma and Drude prescriptions for NbTiN is on the order of $0.2~\mathrm{mPa}$ [14]. Across a narrow sweep just above and below $T_{C}$, conventional models span outcomes ranging from no detectable change to shifts of a few milli-Pascals, depending on how superconducting electrodynamics enters the zero-frequency contribution.
Although superconductivity primarily modifies the optical response at low frequencies, finite-temperature Casimir forces are uniquely sensitive to this regime through the Matsubara zero-frequency (n=0) term. We therefore introduce a model-agnostic approach, analyzing all plausible combinations of Drude, plasma, and BCS permittivity models above and below the superconducting transition. Nearly all combinations predict only smooth or negligible changes across $T_{C}$; the sole exception is the plasma–BCS case, which predicts an abrupt milli-Pascal-scale jump in the pressure gradient (Table 1). This prediction defines the experimental sensitivity threshold that superconducting Casimir platforms must reach.
@@ Acknowledgements.
To highlight the strength of our fabrication approach, Table S1 compares the mechanical properties of membranes suspended over the two extremes of gap size used in this work. Despite a more than fivefold increase in gap, introduced by a thicker silicon spacer, the devices remain nearly identical in resonance frequency, stress, and density, showing that the release process introduces minimal stiction or distortion. This reproducibility is crucial, since it ensures that any variations in response can be attributed to Casimir forces rather than uncontrolled changes in membrane mechanics. The ability to fabricate such matched devices provides a solid foundation for systematically probing Casimir effects across a wide range of separations.
@@
Quality factor $Q$ ($\sim 4.45$ K) ($\times 10^{3}$) 720 200 - 300
Quality factor $Q$ ($\sim 14.2$ K) ($\times 10^{3}$) 1000 130 - 170
@@
[…]
@@
[…]
@@
    Cryogenic SF6 Undercut: The final undercut is performed at –120 C using SF6 plasma. Prior to etching, the sample is cooled down in a high-vacuum environment with a 10 sccm helium flow for 15 min. The etch is then conducted for 3 min and 10 s with SF6 plasma, where the RF power is set to 0 W and the ICP power to 1000 W. After etching, the sample is warmed to 40 C, pumped out, and then unloaded from the process chamber.
@@
    High-Vacuum Baking: The STM holder is placed inside the STM load lock and baked at 200 C under high vacuum ($10^{-8}$ mbar) for two days. This step removes contaminants such as native oxides and residual organic materials.
@@
    Tip Sharpening on an Atomically Flat Copper Sample: Prior to measuring the Casimir samples, an atomically flat copper sample is used to pre-condition the STM tip. This involves: (1) Pulsing the tip with appropriate voltages, and (2) Dipping the tip into the copper sample. These steps refine the tip’s profile, yielding atomic resolution as confirmed by imaging a step height portfolio on the copper sample, as shown in Figure S10(b).
@@
[…]
@@
    Temperature Sweep Considerations: During measurements of the membrane’s resonance frequency as a function of temperature, it is important that the STM tip cannot be retracted out of the tunneling range in the $z$-direction, since the piezo-actuator’s expansion or contraction in the xx/$y$-directions with temperature variations can cause the loss of the original measurement spot. As a result, the atomic tracker, a feedback system in the STM for tracking the same spot on the topography, is activated not only during high-resolution measurement (with a radius of 10 pm) but also during temperature changes (with a tracking radius of approximately 100 pm).
@@
$\Delta\omega^{2}=-\frac{\epsilon_{0}}{\rho h}\frac{(V_{\rm bg}-V_{0})^{2}}{d^{3}}-\frac{\epsilon_{0}}{\rho h}\frac{\delta V_{\rm bg}^{2}}{d^{3}}~,$ (S.1)
where $\epsilon_{0}$ is the vacuum permittivity, $\rho$ is the film density, $h$ is the membrane thickness, $V_{\rm bg}$ is the backgate voltage, and $V_{0}$ is the residual electrostatic voltage. Fitting the experimental data with the above formula, yields the densities ($\rho$) of the suspended NbTiN membranes, while the peak frequencies of the parabolic response curves provide the film stresses.
Supplementary Information C4: Spectroscopy (II-$V$) measurement for checking the bias voltage’s influence on the membrane frequency
Figure S12: (a) Schematic of the spectroscopy (II-$V$) measurement, where the separation between the STM tip and the suspended membrane is kept the same, by varying the bias voltage between the STM and the membrane, and turning off the feedback control of the tunneling current. Tunneling current II as a function of the bias voltage ($V$), between the STM tip and the (b) small- and (c) big-gap membranes, for each pair of II-$V$ values, the tip-sample separation are kept. These values are used for checking whether the tip-sample interaction depends on the bias voltage, as well as the dependence on the temperature, as shown in (d) for small-gap membrane, and in (e) for big-gap membrane.
Experimental results on the nanomembranes’ resonant frequencies dependence on the bias voltage is shown in Figure 3(d) in the main text. The experiment is performed by setting the separation between the STM tip and the membrane unchanged, and varies the bias voltage and thus the tunneling current between the tip and the membrane. The corresponding tunneling current values at different voltages are shown in Figure S12(b) for small-gap and (c) for big-gap membranes. The insets in Figure 3(d) are shown with more details in Figure S12(d) for small-gap and (e) for big-gap membranes
Supplementary Information D: Measurements with atomic force microscopy and white light interferometry
@@
We performed atomic force microscopy (AFM) measurement to identify the roughness of the NbTiN suspended membranes and backgates (bottom films) after the cryogenic SF6 plasma undercut. In Figure S13, we show the AFM scan of the backgate films on the small-gap (Panel (a)) and big-gap (Panel (c)) chips, whose RMS roughness are $(0.90\pm 0.05)$ nm and $(0.96\pm 0.07)$ nm, respectively. The area exposed to the plasma etchant is roughened up to $(1.73\pm 0.55)$ nm and $(1.78\pm 0.80)$ nm for the small- and big-gap bottom films, respectively. The zoom-in images of the bottom films under the small- and big-gap cavities, are shown in Panel (b) and Panel (d), respectively, showing the patch sizes of the NbTiN films to be around $\ell=20\sim 30$ nm in radius.
@@
In Figure S14, we scan the top of the membranes on small- (Panel (a)) and big-gaps (Panel (b)), after they collapsed to the bottom. The roughness of the top membrane over the small-gap cavity is $(1.50\pm 0.27)$ nm, while the one over the big-gap cavity is $(2.73\pm 0.26)$ nm, from which we know that depositing a thicker amorphous silicon layer (1213 nm instead of 190 nm), will increase the roughness of the top NbTiN film a bit.
@@
By comparing the heights of the collapsed membranes over the small-gap (Panel (b)) and big-gap (Panel (d)) cavities in Figure S15, we found both the thicknesses of the membranes over the small- and big-gap to be around 155 nm. From Panels (a) and (c), the small-gap size is found to be (345-155)=190 nm, and the big gap size is found to be (1368-155)=1213 nm, both matching the values we measured with the SEM. Here the suspended films are fabricated with large opening windows on the top NbTiN films, thus large overhangs are realized and the film stresses are released. for the convenience of AFM scan,
@@
Supplementary Information D3: Flatness estimation of the small-gap cavity
In order to estimate the bending amplitude of the NbTiN membrane suspended over the small-gap, we use both theoretical estimation and FEM simulation for the approximation. The static deflection of the membrane due to external forces, such as the electrostatic force and the Casimir force, can be approximated by the formula describing the membrane’s deflection under an uniform pressure $P(z_{0})$ [51]:
$P(z_{0})=C_{1}\frac{4t\sigma}{L^{2}}z_{0}+C_{2}\frac{16E}{L^{4}}z_{0}^{3},$ (S.2)
where $C_{1}$ is a constant and $C_{1}=3.45$ for square membrane, $C_{2}$ is a Poisson-ratio-dependent constant, $E$ is the Young’s modulus, $z_{0}$ is the deflection amplitude, $t$ is the thickness of the membrane, $\sigma$ is the film stress, $L$ is the lateral size of the membrane. For cases of the membranes under high tensile stress, the first term dominates and $P(z_{0})\approx 4C_{1}t\sigma z_{0}/L^{2}$. By comparing it with the static pressure $P_{Cas}(d)$ induced by the Casimir force is
$z_{0,Cas}(d)\approx P_{Cas}(d)\cdot\frac{L^{2}}{4C_{1}t\sigma},$ (S.3)
where $d$ is the gap size. When the square membrane is patterned with holes, an correction factor $C_{hole}$ is required for correcting the deflection:
$z_{0,Cas}(d)\approx C_{hole}P_{Cas}(d)\cdot\frac{L^{2}}{4C_{1}t\sigma}.$ (S.4)
In our case, $C_{hole}\approx 1.086$ (FEM simulation), $P_{Cas}(190~nm)\approx-1.081\times 10^{-24}\times d^{-3.507}$, thus the deflection can be estimated as
$z_{0,Cas}(190~\rm nm)=-1.085\times\frac{1.081\times 10^{-24}}{(190\times 10^{-9})^{3.507}}\times\frac{(709\times 10^{-6})^{2}}{3.45\times 4\times 155*10^{-9}\times 677\times 10^{6}}[m]=-152~pm,$ (S.5)
which is close to the FEM simulation result as shown in Figure S16, whose estimated deflection is 153 pm. For comparison, the big-gap membrane has a maximum deflection under the Casimir pressure $P_{Cas}(1213~nm)\approx-1.013\times 10^{-26}\times d^{-3.829}$, to be $z_{0,Cas}(1213~\rm nm)=0.17~pm$.
Figure S16: Membrane deflection profiles simulated with finite element method. (a) Deflection of the simple $709\times 709\rm~um^{2}$ square membrane, with a maximum deflection of 141 pm. (b) (a) Deflection of the $709\times 709\rm~um^{2}$ square membrane patterned with arrays of holes, with a maximum deflection of 153 pm. A ratio of $C_{hole}=1.085$ is found between the two cases. The FEM simulations are performed by executing a separation-dependent Casimir pressure on the membrane, such that the balance between
Supplementary Information D4: White-light interferometry upper bound on membrane deflection
To place an upper bound on the maximum deflection of the small- and large-gap membranes, we measured their surface profiles using a white-light interferometer (Bruker K1).The measured height maps are shown in Figure S17. Cross-sectional profiles were taken along the blue dashed lines in Figure S17(a,b) and fitted with a cosine function. The fitted amplitudes are $(0.327\pm 0.535)$ nm for the small-gap membrane and $(0.032\pm 0.451)$ nm for the large-gap membrane, both consistent with near-flat profiles within uncertainty. Unlike indirect approaches such as the balanced-capacitance method [7], this direct interferometric measurement provides a conservative upper bound on membrane deflection and supports ultrahigh parallelism between the compliant nanomechanical membrane and the fixed plate. The measurement precision is limited by the interferometric wavelength, diffraction/interference from the membrane hole array, surface roughness, and instrument noise.
Figure S17: White-light interferometry of membrane flatness (Bruker K1). (a,b) Surface-height maps of the small-gap and large-gap membranes. (c,d) Cross-sectional height profiles extracted along the blue dashed lines in (a,b), respectively. Red curves are cosine fits; red shaded bands denote fit uncertainty. Extracted deflection amplitudes are $(0.327\pm 0.535)$ nm (small-gap) and $(0.032\pm 0.451)$ nm (large-gap), both consistent with near-flat profiles within uncertainty. Negative height corresponds to deflection toward the backgate electrode.
@@
$G(V_{\rm bias})=\frac{dI}{dV_{\rm bias}}=A\int^{+\infty}_{-\infty}N(E,\gamma,\Delta)\frac{\partial f(E+e\cdot V_{\rm bias},T)}{\partial(e\cdot V_{\rm bias})}dE,$ (S.6)